### v1

# Criticality Leveraged Adversarial Training (CLAT) for Boosted Performance via Parameter Efficiency

Abstract
Adversarial training enhances neural network robustness but suffers from a tendency to overfit and increased generalization errors on clean data. This work introduces CLAT, an innovative approach that mitigates adversarial overfitting by introducing parameter efficiency into the adversarial training process, improving both clean accuracy and adversarial robustness. Instead of tuning the entire model, CLAT identifies and fine-tunes robustness-critical layers—those predominantly learning non-robust features—while freezing the remaining model to enhance robustness. It employs dynamic critical layer selection to adapt to changes in layer criticality throughout the fine-tuning process. Empirically, CLAT can be applied on top of existing adversarial training methods, significantly reduces the number of trainable parameters by approximately 95%, and achieves more than a 2% improvement in adversarial robustness compared to baseline methods.
1 Introduction
Advancements in deep learning models have markedly improved image classification accuracy. Despite this, their vulnerability to adversarial attacks — subtle modifications to input images that mislead the model — remains a significant concern [9, 28]. The research community has been rigorously exploring theories to comprehend the mechanics behind adversarial attacks [4]. Ilyas et al. [14] uncover the coexistence of robust and non-robust features in standard datasets. Adversarial vulnerability largely stems from the presence of non-robust features in models trained on standard datasets, which, while highly predictive and beneficial for clean accuracy, are susceptible to noise [28]. Unfortunately, it is observed that deep learning models tend to preferentially learn these non-robust layers. Inkawhich et al. [15, 16] further demonstrate that adversarial images derived from the hidden features of certain intermediate non-robust/“critical” layers exhibit enhanced transferability to unseen models. This suggests a commonality in the non-robust features captured by these layers. While identifying these critical layers to improve their robustness is appealing, this process often requires the costly generation of attacks against each individual network layer. Efficient methods to identify and effectively address the criticality of such layers are still lacking.
In contrast to layer-wise feature vulnerability analysis, adversarial training [3, 19, 8], involves training entire neural networks with adversarial examples generated in real-time. This approach inherently encourages all layers in the model to learn robust features from adversarial images, thereby enhancing the model’s resilience against attacks. However, given the more challenging optimization process of learning from adversarial examples than from clean examples, adversarial training also brings hurdles such as heightened errors on clean data and susceptibility to overfitting, ultimately diminishing its effectiveness in practical applications [25, 36, 23, 17]. Despite various efforts to enhance adversarial training, such as modifying input data and adjusting loss functions [12, 23, 36, 30, 32, 21], they still frequently fall short in alleviating the aforementioned challenges.
Figure 1: CLAT overview. CLAT fine-tunes the selected critical layers (red) while freezing other layers (grey). fine-tuning objective is computed per Equation 9. Critical layers are adjusted periodically. Pseudocode is provided in Appendix A.
In light of these challenges, we introduce CLAT, a paradigm shift in adversarial training, where we mitigate overfitting during adversarial training by identifying and tuning only the robustness-critical model layers. CLAT commences by pinpointing critical layers within a model using our novel, theoretically grounded, and easily computable,“criticality index”, which we developed to identify layers which have learned non-robust features dominantly. Subsequently, our algorithm meticulously fine-tunes these critical layers to remove their non-robust features and reduce their criticality, while freezing the other, non-critical layers. Dynamic selection of critical layers is conducted during the training process to always focus the fine-tuning on the most-in-need layers, avoiding the overfitting of full-model adversarial training. CLAT therefore enhances both clean and adversarial accuracy compared to previous adversarial training methods. Besides reducing overfitting, the critical layer selection in training also makes CLAT a streamlined, parameter-efficient fine-tuning method, enabling memory savings with its significantly reduced trainable parameters.
In summary, we make the following contributions in this paper:
  • We introduce the “criticality index”, a quantitative metric designed to efficiently identify critical layers for the adversarial vulnerability of the model.
  • We develop a specialized adversarial training objective focused on reducing the criticality of the identified critical layers to bolster overall model robustness.
  • We propose CLAT, a parameter-efficient adversarial fine-tuning algorithm focusing on layer criticality reduction. CLAT mitigates overfitting in the adversarial training process, and applies to various model training scenarios and baseline adversarial training methods.
CLAT stands out by markedly reducing overfitting risks, maintaining or even slightly improving the clean accuracy while significantly enhancing adversarial resilience by at least 2% with a 96% reduction in trainable parameters.
2 Related Work
2.1 Adversarial training
Adversarial Training (AT) was first introduced by Goodfellow et al. [9], who demonstrated how the integration of adversarial examples into the training process could substantially improve model robustness. This idea evolved into a sophisticated minimax optimization approach with Projected Gradient Descent Adversarial Training (PGD-AT) [19], which employs PGD attacks in training. Regarded as the gold standard in AT, PGD-AT generates adversarial training samples using multiple steps of projected gradient descent, leading to substantially improved empirical robustness [6, 3, 8]. Further refining this approach, TRADES [36] optimizes a novel loss function to balance classification accuracy with adversarial robustness. Recent enhancements in AT, including model ensembling and data augmentation, have also produced notable improvements in model resilience. [33, 34, 7]Inkawhich et al. [15] propose “Activation Attacks” (AA) which underscore the efficacy of leveraging intermediate model layers for generating stronger adversarial attacks, suggesting that incorporating AA in adversarial training could fortify defenses. Their findings provide a foundation for our method which integrates these intermediate critical layers into our adversarial training strategy.
2.2 Efficient adversarial training
Adversarial training methods like PGD-AT and TRADES are computationally intensive, necessitating the multi-step generation of adversarial examples, optimization of complex objective functions, and comprehensive model tuning [26]. Numerous efforts to enhance the efficiency of adversarial training have been proposed. Shafahi et al. [26] introduced an efficient model known as “Free” AT, using a single backpropagation step for both training and PGD adversary generation to accelerate adversarial training. However, this approach struggled with maintaining model robustness and increased overfitting, primarily due to gradient alignment issues. Wong et al. [31] then proposed “Fast” AT, which aimed to address these shortcomings, but was later found to be vulnerable to similar problems [2]. As a result, Andriushchenko and Flammarion [2] developed GradAlign but this approach tripled the training time due to second-order gradient computation. Lastly, RiFT leverages network layer redundancies to improve general performance, but its heuristic-based layer redundancy measurement limits final robustness [37]. In contrast, CLAT uses a theoretically grounded, dynamic, critical layer selection mechanism, resulting in a more robust and parameter-efficient adversarial training algorithm. Furthermore, CLAT does not require a fully trained model and can be combined with existing fast-AT methods to enhance performance and mitigate overfitting.
3 Methods
Building on prior attack and defense research [15, 16, 37] which demonstrates that not all model layers equally learn non-robust features, we aim to improve model robustness and parameter efficiency by identifying and fine-tuning only those critical layers that are prone to learning non-robust features, while keeping the non-critical layers frozen. In this section, we begin by defining and identifying critical layers, then outline our objectives for reducing their criticality. Finally, we present our complete CLAT algorithm, which effectively enhances model robustness.
3.1 Layer Criticality
Consider a deep learning model with $n$ layers, and an input $x$, defined as:
$F(x)=f_{n}(f_{n-1}(\dots f_{1}(x))),$ (1)
where the functionality of the $i$-th layer is denoted as $f_{i}$. During the standard training process, all layers learn useful features which contribute to the correct outputs of the model. We denote the hidden feature learned at the output of the $i$-th layer as $F_{i}(x)=f_{i}(f_{i-1}(\dots f_{1}(x)))$.
Under adversarial perturbation, features from all layers will be altered, leading to incorrect outputs. To understand how the learned parameters of the model affect the features under perturbation, we consider a function $G_{i}(\cdot)$ defined at the output of layer $i$ on inputs close to a clean data point $x$:
$G_{i}(z)=||F_{i}(z)-F_{i}(x)||_{2}^{2}.$ (2)
Clearly, $z=x$ is a minima of $G_{i}(z)$. The worst-case curvature of the function $G_{i}$ at the neighborhood of $z=x$ can be estimated following the formulation by Moosavi-Dezfooli et al. [20] as
$\nu_{i}(x)=\frac{\nabla G_{i}(x^{\prime})-\nabla G_{i}(x)}{||x^{\prime}-x||_{2}}=\frac{\nabla G_{i}(x^{\prime})}{||x^{\prime}-x||_{2}},$ (3)
where $x^{\prime}$ is a worst-case perturbation (adversarial attack) maximizing $G_{i}(z)$ in the vicinity of $x$, and $\nabla G_{i}(x)=0$ by definition given it is a minimum.
The numerator in the curvature formulation can be further derived as
$\nu_{i}(x)=\frac{\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}^{T}(F_{i}(x^{\prime})-F_{i}(x))}{||x^{\prime}-x||_{2}}.$ (4)
Following the observation in [20], a higher curvature indicates the feature to be more sensitive to adversarial examples. We can therefore use the curvature formulation $\nu_{i}(x)$ under a fixed perturbation budget $||x^{\prime}-x||_{p}\leq\epsilon$ to estimate the layer sensitivity, or weakness.
In practice, it is difficult to explicitly instantiate $\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}$ for a neural network. To this end, we simplify the formulation in Equation 4 by assuming $\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}$ as a uniform vector. This leads to our definition of the $\epsilon$-weakness of layer $i$’s feature as:
$\mathcal{W}_{\epsilon}(F_{i})=\frac{1}{N_{i}}\mathbb{E}_{x}\left[\sup_{||\delta||_{p}\leq\epsilon}||F_{i}(x+\delta)-F_{i}(x)||_{2}\right],$ (5)
where $N_{i}$ denotes the dimensionality of the output features at layer $i$, therefore normalizing the weakness measurement of layers with different output sizes. The weakness measurement is proportional to the curvature estimation in Equation 4. A higher weakness value indicates that the feature vector is more vulnerable to input perturbations. The functionality of cascading layers from 1 to $i$ affects the vulnerability of the hidden features, as described by this formulation.
For the purpose of efficient fine-tuning, we want to identify the layers that are the most critical to the lack of robustness, characterized by their increased susceptibility to perturbations from adversarial inputs. We therefore provide the following definition:
Definition 3.1.
Critical layer: A layer is considered critical if it exhibits a greater propensity to learn non-robust features or demonstrates heightened sensitivity to adversarial input perturbations relative to other layers in the model.
To this end, we single out the contribution of each layer’s functionality to the weakness of the features after it with a Layer Criticality Index $\mathcal{C}_{f_{i}}$, which is formulated as
$\mathcal{C}_{f_{i}}=\frac{\mathcal{W}_{\epsilon}(F_{i})}{\mathcal{W}_{\epsilon}(F_{i-1})}.$ (6)
For the first layer, we define $\mathcal{C}_{f_{1}}=\mathcal{W}_{\epsilon}(F_{1})$ as only the first layer contributes to the weakness.
As a sanity check, the feature weakness at the output of layer $i$ can be attributed to the criticality of all previous layers as $\mathcal{W}_{\epsilon}(F_{i})=\prod_{k=1}^{i}\mathcal{C}_{f_{k}}$. Conversely, a layer with a larger criticality index will increase the weakness of the features after it, indicating the layer is critical according to Definition 3.1, as it heightens the feature sensitivity to the adversarial input.
One drawback of the formulation in Equation (6) is that computing the feature weakness involves finding the worst-case perturbation against the hidden features at each layer, which is a costly process to conduct sequentially for all layers. In practice, we find it possible to approximate the worst-case perturbation against features with an untargeted PGD attack against the model output, so that we can use the same PGD perturbation $\delta$ to estimate the feature weakness of all layers following Equation 5. In this way, with a reasonably sufficient batch size, we can compute the critical indices for all layers in a model with two forward passes: one with the clean input $x$ and one with the PGD attack input $x+\delta$. We make the following proposition:
Proposition 3.2.
Critical layers defined as in Definition 3.1 can be identified as the layers with the largest criticality indices $\arg\max_{i}\mathcal{C}_{f_{i}}$.
To verify Proposition 3.2, we conduct an ablation study in Table 6, where we show that model robustness is improved more by CLAT fine-tuning of critical layers compared to equivalent fine-tuning of randomly selected layers. We will discuss how to reduce the criticality of the critical layers and make them more robust in the next subsection.
3.2 Criticality-targeted fine-tuning
Once the critical layers are identified, we fine-tune them to reduce their criticality, thereby decreasing the weakness of subsequent hidden features and enhancing model robustness. For a critical layer $i$, we optimize the trainable parameters to minimize $\mathcal{C}_{f_{i}}$. Note that in the criticality formulation in Equation 6, the weakness of the previous layer’s output, $\mathcal{W}_{\epsilon}(F_{i-1})$, is constant with respect to $f_{i}$. Thus, the optimization objective for $f_{i}$ can be simplified as
$\mathcal{L}_{C}(f_{i})=\mathbb{E}_{x}\left[\sup_{||\delta||_{p}\leq\epsilon}||F_{i}(x+\delta)-F_{i}(x)||_{2}\right].$ (7)
In the case where multiple critical layers are considered in the fine-tuning process, the fine-tuning objective can be expanded to accommodate all critical layers simultaneously. Formally, suppose we have a set $S$ where layers $i\in S$ are all selected for fine-tuning, the fine-tuning objective for these critical layers can be formulated as
$\mathcal{L}_{C}(f_{S})=\mathbb{E}_{x}\left[\sup_{||\delta||_{p}\leq\epsilon}\sum_{i\in S}||F_{i}(x+\delta)-F_{i}(x)||_{2}\right],$ (8)
where a single perturbation is utilized to capture the weakness across all critical layers. In practice, a projected gradient ascent optimization with random start is used to solve the inner maximization.
Minimizing the objective in Equation 8 by adjusting the trainable variables of the critical layers will reduce their feature weaknesses. However, the removal of non-robust features in these layers may affect the functionality of the model on clean inputs. As a tradeoff, we also include the cross entropy loss $\mathcal{L}(\cdot)$ in the final optimization objective, which derives the optimization objective on the critical layers during the fine-tuning process
$\min_{f_{S}}\mathbb{E}_{x,y}\mathcal{L}(F(x),y)+\lambda\mathcal{L}_{C}(f_{S}),$ (9)
where the hyperparameter $\lambda$ serves as a balancing factor between the two loss terms. Note that only the selected critical layers $f_{S}$ are optimized in Equation 9 while the other non-critical layers are frozen, justifying our fine-tuning to be parameter-efficient.
3.3 CLAT adversarial training
Similar to parameter-efficient fine-tuning techniques in other domains, we design CLAT as a fine-tuning approach, which is applied to neural networks that have undergone preliminary training. The pretraining phase allows all layers in the model to capture useful features, which will facilitate the identification of critical layers in the model. Notably, CLAT’s versatility allows it to adapt to various types of pretrained models, either adversarially trained or trained on a clean dataset only. In practice, we find that models do not need to fully converge during the pretraining phase to benefit from CLAT fine-tuning. For example, in case of the CIFAR-10 dataset, 50 epochs of PGD-AT training would be adequate. We consider the number of adversarial pretraining epochs as a hyperparameter and provide further analysis on the impact of pretraining epochs in Section 4.3.1.
After the pretraining, CLAT begins by identifying and selecting critical layers in the pretrained model. We then fine-tune critical layers only while freezing the rest of the layers. This process is illustrated in Figure 1, and the pseudocode is provided in Algorithm 1 in Appendix A for greater clarity. Compared to full-model fine-tuning, CLAT not only improves parameter efficiency with fewer trainable parameters, but also reduces overfitting, a common issue in standard adversarial training. As fine-tuning progresses, the critical layers will be updated to reduce their criticality, making them less critical than some of the previously frozen layers. Subsequently, we perform periodic reevaluation of the top $k$ critical indices, ensuring continuous adaptation and optimization of the layers that are the most in need in the training process. Through hyperparameter optimization, we find 10 epochs to be adequate to optimize the selected critical layers for all models that we tested.
4 Experiments
4.1 Experimental settings
Datasets and models
We conducted experiments using two widely recognized image classification datasets, CIFAR10 and CIFAR100. Each dataset includes 60,000 color images, each 32×32 pixels, divided into 10 and 100 classes respectively [18]. For our experiments, we deployed a suite of network architectures: WideResnets (34-10, 70-16) [35], ResNets (50, 18) [11], DenseNet-121 [13], PreAct ResNet-18[10], and VGG-19 [27]. In this paper, these architectures are referred to as WRN34-10, WRN70-16, RN50, RN18, DN121, Preact RN18 and VGG19 respectively.
Training and evaluation
Since CLAT can be layered over clean pretraining, partial training, or other adversarial methods, results incorporating CLAT are denoted in our tables as "X+ CLAT," where “X” specifies the baseline method used prior to applying CLAT. Typically, this baseline method was applied for the first 50 epochs, followed by a fine-tuning phase during which CLAT was run for an additional 50 epochs. Models trained exclusively with PGD-AT typically require 150 epochs to achieve reported performances. We use PGD for attack generation unless stated otherwise in the baseline method, which employs a random start [19], with an attack budget of $\epsilon$ = 0.03 under the $\ell_{\infty}$ norm, an update step size of $\alpha$ = 0.007, and 10 attack steps. Experiments were conducted on a Titan XP GPU, starting with an initial learning rate of 0.1, which was adjusted according to a cosine decay schedule. To ensure the reliability of robustness measurements, we conducted each experiment a minimum of 10 times, reporting the lowest adversarial accuracies we observed.
CLAT settings
We select critical layers as described in Section 3.1. Table 4 outlines the Top-5 most critical layers for some of the models and corresponding datasets at the start of the CLAT fine-tuning, after adversarially training with PGD-AT for 50 epochs. In customizing the CLAT methodology to various network sizes, we select approximately 5% of layers as critical through hyperparameter optimization. For instance, DN121 uses 5 critical layers, while WRN70-16, RN50, WRN34-10, VGG19, and RN18 use 4, 3, 2, 1, and 1 critical layers, respectively.
4.2 CLAT performance
White-box robustness
Table 1 presents the clean accuracy and white-box $\ell_{\infty}$ PGD [6] accuracy across various adversarial training techniques. The versatility and effectiveness of CLAT are demonstrated by combining it with various standard adversarial training methods. CLAT mitigates the overfitting commonly observed in traditional adversarial training methods, improving both clean and adversarial accuracy compared to the corresponding baselines.
We also show that reducing trainable parameters alone does not guarantee a performance improvement by comparing CLAT against other parameter-efficient fine-tuning methods like LoRA [1] and RiFT [37]. Our theoretically-grounded criticality indices enable the precise identification of critical layers and the focused elimination of non-robust features from these pivotal layers. Furthermore, we verify that fast adversarial training techniques as in [31] can be applied to solve the inner maximization in the CLAT training objective in Equation 8, where the “CLAT (Fast)” method also improves performance and robustness over Fast-AT baselines.
Besides the main table, we provide additional robustness results in Appendix B, where we showcase CLAT’s improved robustness against a variant of the white-box Auto Attack [8] in LABEL:wb_aa and verify the generalizability of the robustness to attacks of various strengths in Figure 4.
Table 1: Comparative performance of CLAT across various networks and adversarial training/fine-tuning techniques. Robustness is evaluated with white-box PGD attacks.
MODEL METHOD CIFAR-10 ACC. (%) CIFAR-100 ACC. (%)
clean adversarial clean adversarial
DN121 pgd-at [19] 80.05 58.15 57.18 31.76
pgd-at + clat 81.03 60.60 58.79 33.23
WRN70-16 Peng et al. [22] 93.27 71.07 70.20 42.61
peng et al. + clat 93.56 72.25 71.94 44.12
Bai et al. [5] 92.23 64.55 69.17 40.86
bai et al. + clat 92.77 64.92 70.17 41.64
RN50 pgd-at 81.38 56.35 58.16 33.01
pgd-at + clat 83.78 59.54 61.88 36.23
WRN34-10 pgd-at 87.41 55.40 59.19 31.66
pgd-at + lora [1] 73.36 56.17 55.56 31.43
pgd-at + rift [37] 87.89 55.41 62.35 31.64
pgd-at + clat 88.97 57.11 62.38 32.05
trades [36] 87.60 56.61 60.56 31.85
trades + rift 87.55 56.72 61.01 32.03
trades + clat 88.23 57.89 61.45 33.56
VGG19 pgd-at 78.38 50.35 50.16 26.54
pgd-at + clat 79.88 52.54 50.98 28.41
RN18 pgd-at 81.46 53.63 57.10 30.15
pgd-at + lora 76.57 55.38 48.49 32.36
pgd-at + rift 83.44 53.65 58.74 30.17
pgd-at + clat 83.89 55.37 59.22 32.04
trades 81.54 53.31 57.44 30.20
trades + rift 81.87 53.30 57.78 30.22
trades + clat 81.89 54.57 58.82 31.06
mart [29] 76.77 56.90 51.46 31.47
mart + rift 77.14 56.92 52.42 31.48
mart + clat 76.82 57.55 53.01 33.23
awp [32] 78.40 53.83 52.85 31.00
awp + rift 78.79 53.84 54.89 31.05
awp + clat 79.01 55.27 55.39 32.08
score [21] 84.20 54.59 54.83 29.49
score + rift 85.65 54.62 57.63 29.50
score + clat 86.11 55.78 57.66 30.23
PREACT RN18 fast-at [31] 81.46 45.55 50.10 27.72
fast-at + clat 84.46 52.13 54.33 29.22
fast-at + clat (fast) 82.72 49.62 52.10 27.99
Black-box robustness
Table 2 evaluates the robustness against black-box attacks between models trained solely using PGD-AT and those augmented with CLAT. Attack settings are the same as those of the white-box attacks. As a sanity check, the accuracies under black-box attack surpass those observed under white-box scenarios, indicating that gradient masking does not appear in the CLAT model, and that the white-box robustness evaluation is valid. More significantly, models trained with CLAT consistently outperform those trained with PGD-AT, maintaining superior resilience in both black-box and white-box settings, regardless of the attack method or models employed.
Table 2: Comparative Analysis of Black-box PGD Accuracy on CIFAR-10 and CIFAR100. Model in each row is the attacker and each column the victim.
Network Method Cifar-10 Adv. Acc. (%) Cifar-100 Adv. Acc. (%)
Rn50 Dn121 Vgg19 Rn18 Rn50 Dn121 Vgg19 Rn18
Rn50 pgd-at - 74.83 68.01 67.44 - 46.82 40.55 40.10
clat - 76.45 71.25 70.12 - 48.49 44.34 43.91
Dn121 pgd-at 72.24 - 69.53 68.38 44.45 - 40.63 41.22
clat 74.55 - 71.78 70.56 46.78 - 43.62 42.88
Vgg19 pgd-at 65.72 67.56 - 62.26 47.86 46.56 - 40.55
clat 66.46 70.72 - 65.78 49.25 48.72 - 42.72
Rn18 pgd-at 74.82 70.21 61.83 - 46.28 45.59 39.21 -
clat 76.23 73.19 63.96 - 48.89 47.72 41.78 -
4.3 Ablation Studies
4.3.1 Ablating on pretraining epochs before CLAT
As discussed in Section 3.3, we apply CLAT after the model has been adversarially trained for some epochs. Here, we analyze how the number of pretraining epochs affects CLAT performance. Figure 2 shows the training curves for different allocations of PGD pretraining epochs and CLAT fine-tuning epochs within a 100-epoch training budget. The overfitting of PGD-AT is evident as adversarial accuracy plateaus and declines towards the end, as documented in previous research [24]. In contrast, CLAT continues to improve adversarial accuracy, effectively addressing this issue. Including CLAT at any stage of training results in higher clean accuracy and robustness at convergence. Additional results on pretrained clean models are provided in Appendix C.
Furthermore, an intriguing aspect of our experiments involves running CLAT from scratch (0 PGD-AT epochs). Although CLAT ultimately surpasses PGD-AT with sufficient epochs, using CLAT without any prior adversarial training results in significantly slower model convergence. We believe this suggests that “layer criticality” emerges during the adversarial training process, allowing critical layers to be identified as the model undergoes adversarial training. This phenomenon supports our theoretical insight that criticality can be linked to the curvature of the local minima to which each layer converges during adversarial training. Flattening the local minima of critical layers aids in model generalization. We plan to explore this phenomenon further in future work.
Figure 2: Comparative analysis of CLAT performance on WRN34-10: Clean and adversarial accuracy on CIFAR-10 across partially trained models. Note: PGD-AT requires 150 epochs to fully converge.
Table 3: Top-5 criticality indices by model and dataset. Layers used in CLAT are bolded.
Model CIFAR10 CIFAR100
DN121 39, 14, 1, 3, 88 39, 15, 1, 2, 91
WRN70-16 4, 17, 1, 59, 62 3, 17, 2, 59, 61
RN50 34, 41, 48, 3, 36 34, 43, 45, 6, 32
WRN34-10 26, 1 30, 3, 28 26, 2, 30, 3, 27
VGG19 9, 11, 5, 3, 1 8, 13, 5, 3, 1
RN18 11, 10, 4, 2, 12 12, 9, 5, 2, 13
Table 4: Comparison of PGD-AT, PGD-AT + CLAT, and PGD-AT + CLAT (Non-dynamic) on CIFAR-10 Adv. Accuracy.
Method DN121 RN50
pgd-at 58.15 56.35
clat 60.60 59.54
clat (nd) 57.01 54.22
4.3.2 Ablating on critical layer selection
The choice of critical layer selection is another important feature impacting the performance of CLAT. We begin by examining the effect of dynamic layer selection. Table 4 highlights that dynamic selection is crucial to CLAT’s performance. Using the same layers throughout the process tends to cause overfitting and results in lower accuracies compared to the PGD-AT baseline.
To verify the significance of the selected critical layers, we compare CLAT with an alternative approach in which random layers are dynamically selected for fine-tuning instead of the critical layers. The results of this comparison are detailed in Table 6 and Table 10 (in Appendix C) on CIFAR-10 and CIFAR-100, respectively. The data demonstrates that selecting critical layers significantly enhances the model’s adversarial robustness and yields a marginal increase in clean accuracy. Moreover, Table 4 indicates near-identical critical layer selections within the same model even across diverse datasets. This evidence supports our assertion that the variation in layer criticality stems from inherent properties within the model architecture, wherein certain layers are predisposed to learning non-robust features.
Table 5: Critical vs. random layers for CLAT on CIFAR-10.
Network CIFAR-10
Critical Layers Random Layers
Clean Acc. Adv. Acc. Clean Acc. Adv. Acc.
DN121 81.03 60.60 78.85 51.35
RN50 83.78 59.54 79.01 51.44
RN18 83.89 55.37 78.02 51.03
Table 6: Trainable Parameters during CLAT in Various Networks
Network Trainable Params % Used
Total clat
dn121 6.96M 217K 3.1%
wrn70-16 267M 8.29M 3.0%
rn50 23.7M 823K 3.4%
wrn34-10 46.16M 1.24M 2.7%
rn18 11.2M 590K 5.2%
vgg19 39.3M 236K 6.0%
Lastly, we ablate on the number of critical layers used in CLAT for fine-tuning. Figure 3 illustrates the trade-off between adversarial test accuracy and the quantity of critical layers selected for fine-tuning. Allowing more layers to be fine-tuned enhances the model flexibility, which initially improves CLAT performance. However, fine-tuning an excessive proportion of critical layers diminishes CLAT’s effectiveness. This is likely due to the diversion of attention towards fine-tuning less-critical layers, detracting from more critical ones. This pattern highlights the crucial influence of specific layers on network robustness and suggests a need for deeper research to understand the roles and dynamics of individual layers influencing network robustness.
Figure 3: Comparative analysis on CLAT performance/adversarial accuracy with respect to number of critical layers used during CLAT
4.4 Efficiency and stability analysis
To verify the significance of the selected critical layers, we contrast CLAT with an alternative approach in which random layers are dynamically selected for fine-tuning instead of the critical layers. The results of this comparison are detailed in Table 6 and Table 10 (in Appendix C) on CIFAR-10 and CIFAR-100, respectively.
The cost of optimizing the CLAT training objective in Equation 9 is similar to that of the standard adversarial training given its min-max formulation. Here, we show that the computational overhead for determining critical indices is negligible. We verify in Table 7 that criticality indices can be stably computed with a single training batch as small as 10, with top-ranking layers consistent with those achieved with a larger batch. We further conducted over 1000 runs for each network to randomly select the data used for criticality estimation, where we find remarkable consistency in computed critical layers, differing in less than 5% of cases, typically involving only one layer change among the top five critical layers. The stability means we can use a mere 0.0002% of the training data for criticality estimation every 10 epochs, which introduces neglectable additional complexity to the training process as indicated by the time measurement in the table.
Table 7: DenseNet-121 critical layers identified with different amount of data. Time taken to compute critical layers evaluated on TITAN RTX GPU. As a reference, 1 PGD-AT epoch takes 67s.
Batch Size Cifar10 Cifar100
Critical Layers Time (s) Critical Layers Time(s)
10 39, 14, 1, 3, 90 2.64 39, 15, 1, 2, 91 2.82
30 39, 14, 1, 3, 88 2.72 39, 15, 1, 2, 88 2.91
50 39, 14, 1, 3, 89 2.83 39, 15, 1, 3, 91 3.14
100 39, 14, 1, 3, 88 3.15 39, 15, 1, 2, 91 3.54
5 Conclusions
In this work, we introduce CLAT, an innovative adversarial training approach that addresses adversarial overfitting issues by fine-tuning only the critical layers vulnerable to adversarial perturbations. This method not only emphasizes layer-specific interventions for enhanced network robustness but also sheds light on the commonality in non-robust features captured by these layers, offering a targeted and effective defense strategy. Our results reveal that CLAT reduces trainable parameters by about 95% while ensuring significant improvements in clean accuracy and adversarial robustness across diverse network architectures and baseline adversarial training methods. We limit the scope of this work to improving the empirical robustness of the model by utilizing the critical layers. In this sense, open questions remain on why these specific layers become critical, whether they can be identified more effectively, and whether the issues can be resolved with architectural or training scheme changes. We leave a more theoretical understanding of these questions as future work.
Appendix A Pseudocode of CLAT
To better facilitate an understanding of the CLAT process, we illustrate the pseudocode of the dynamic critical layer identification process and the criticality-targeted fine-tuning process in Algorithm 1. Only the selected critical layers are being fine-tuned while all the other layers are frozen.
Algorithm 1 CLAT Algorithm
1:Input: Dataset $\mathcal{D}$, pre-trained model $F$, batch size $bs$, total epochs $N$.
2:for $epoch=1$ to $N$ do
3:     if $epoch\%10==1$ then
4:         # Find critical layers
5:         $x\leftarrow$ Batch of training data in $\mathcal{D}$
6:         $x+\delta\leftarrow$ PGD attack against $F$
7:         Compute $\mathcal{W}_{\epsilon}(F_{i})$ for all layers with Equation 5
8:         Compute $\mathcal{C}_{f_{i}}$ for all layers with Equation 6
9:         Critical layers $\mathcal{S}\leftarrow TopK(\mathcal{C}_{f_{i}})$
10:     # fine-tune critical layers
11:     $minibatches\leftarrow CreateMinibatches(\mathcal{D},bs)$
12:     for $x,y$ in $minibatches$ do
13:         Perturbation $\delta\leftarrow$ Equation 8 inner maximization
14:         $\mathcal{L}_{C}(f_{\mathcal{S}})\leftarrow$ Equation 8
15:         Weight update $w[\mathcal{S}]$ with Equation 9
Appendix B Additional experiment results
Table LABEL:wb_aa showcases results under a variant of the white-box Auto Attack [8] with an epsilon value of 0.03 and a single restart in Table 1. Specifically, we report adversarial accuracies of the “APGDCE” untargeted attack, a component of the Auto Attack ensemble, with similar trends observed for other attacks in the ensemble.
Although CLAT models employ PGD-like attacks on hidden features during the adversarial training phase and have not seen Auto Attacks, the robustness remains under Auto Attacks. This highlights the similarity in non-robust features utilized by different attacks in different networks. As CLAT resolves the non-robust features with critical layer fine-tuning, the resulting robustness is generalizable across attack settings and source models of the attacks.
Table 8: Adversarial accuracy under the white-box “APGDCE” variant of Auto Attack: Models trained using CLAT consistently outperform those trained exclusively with PGD-AT.
Dataset Network PGD-AT(%) PGD-AT + CLAT (%)
CIFAR-10 DN121 49.26 50.58
WRN70-16 54.48 56.90
RN50 48.54 50.65
WRN34-10 52.21 54.73
VGG19 45.35 47.01
RN18 46.99 48.20
CIFAR-100 DN121 24.67 26.31
WRN70-16 30.27 33.32
RN50 25.44 26.64
WRN34-10 27.62 31.53
VGG19 22.50 24.02
RN18 23.81 24.43
We further compare the CLAT model robustness with the robustness of the SAT model against white-box attacks of various strengths. As illustrated in  Figure 4, though both models are only trained against an attack of one strength ($\epsilon$ = 0.03), the improved robustness of CLAT is consistent across the full spectrum of attack strengths. This shows that CLAT is not overfitting to the specific attack strength used in training.
Figure 4: White-box adversarial accuracy (y-axis) on CIFAR-10 for models trained with CLAT (red) and PGD-AT (blue), against PGD attacks of varying strengths (x-axis)
Appendix C Additional ablation study
C.1 CLAT on pretrained clean model
Besides the discussion on performing CLAT after adversarial pretraining in Section 4.3.1, Table 9 details the performance of CLAT on clean pretrained models. Although the adversarial accuracies of clean pretrained models are relatively low compared to those of adversarially trained models, CLAT demonstrates its capability to facilitate adversarial fine-tuning on clean models effectively to some extent. This is a novel achievement, showcasing the algorithm’s versatility.
Table 9: Adversarial and clean accuracies for performing CLAT on various PyTorch pretrained models on the CIFAR-10 dataset.
Model Adv. Acc. Clean Acc.
dn-121 39.21% 80.89%
wrn70-16 42.1% 83.35%
rn-50 35.67% 78.23%
wrn34-10 40.1% 81.78%
vgg-19 32.67% 75.05%
rn-18 34.45% 76.51%
C.2 Additional results on layer selection
Here in Table 10, we provide additional results contrasting CLAT with an alternative approach where random layers are dynamically selected for fine-tuning instead of the critical layers on the CIFAR-100 dataset.
Table 10: Ablating the layer choices for CLAT fine-tuning on CIFAR-100.
Network CIFAR-100
Critical Layers Random Layers
Clean Acc. Adv. Acc. Clean Acc. Adv. Acc.
DN121 58.79 33.23 54.85 23.32
RN50 61.88 36.23 56.50 22.45
RN18 59.22 32.04 52.43 22.60
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Main claims made in the abstract and introduction accurately reflect the scope and contribution of the paper, supported by our theoretical derivations and experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: This paper is limited to proposing an empirically effective method to identify adversarially critical layers and reduce their criticality, without answering the question on why such discrepancy in criticality appears in the model in the first place. The limitation itself and potential future directions to resolve it are discussed in Section 5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Instead of providing theoretical proof, we justify the formulation of our proposed method empirically through a set of comparisons against baselines (Section 4.2) and detailed ablation studies (Section 4.3).
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We discuss the experiment settings in Section 4.1.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All experiments are done with publicly-available datasets. Besides discussing our settings in Section 4.1, we will release the code and model checkpoints upon the acceptance of the paper.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See Section 4.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We train all of our models with a randomly selected fixed random seed. For evaluation, we conduct adversarial attacks on each model for at least 10 times, and report the lowest adversarial accuracy (instead of an error bar) as it reflects the true robustness of the model more reliably.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report the computation equipment used in Section 4.1. We further provide detailed discussion in Section 4.4 on the difference in cost between our proposed method and previous standard adversarial training research.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We confirm that our paper, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: This paper provides a method that improves the robustness of deep neural networks against adversarial attacks, which helps promoting trustworthy deep learning applications. No potential misuse or negative social impact can be think of.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No potential misuse or negative social impact can be think of.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite all the assets we used properly.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We will provide code and pretrained models with proper documentations.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We have no human subjects in our paper.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We have no human subjects in our paper.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 RN18 59.22 32.04 52.43 22.60
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: Main claims made in the abstract and introduction accurately reflect the scope and contribution of the paper, supported by our theoretical derivations and experimental results.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9.
-Justification: This paper is limited to proposing an empirically effective method to identify adversarially critical layers and reduce their criticality, without answering the question on why such discrepancy in criticality appears in the model in the first place.
-The limitation itself and potential future directions to resolve it are discussed in Section 5.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14.
-Justification: Instead of providing theoretical proof, we justify the formulation of our proposed method empirically through a set of comparisons against baselines (Section 4.2) and detailed ablation studies (Section 4.3).
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: We discuss the experiment settings in Section 4.1.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24.
-Justification: All experiments are done with publicly-available datasets.
-Besides discussing our settings in Section 4.1, we will release the code and model checkpoints upon the acceptance of the paper.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: See Section 4.1.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No]
-  34.
-Justification: We train all of our models with a randomly selected fixed random seed.
-For evaluation, we conduct adversarial attacks on each model for at least 10 times, and report the lowest adversarial accuracy (instead of an error bar) as it reflects the true robustness of the model more reliably.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39.
-Justification: We report the computation equipment used in Section 4.1.
-We further provide detailed discussion in Section 4.4 on the difference in cost between our proposed method and previous standard adversarial training research.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We confirm that our paper, in every respect, with the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49.
-Justification: This paper provides a method that improves the robustness of deep neural networks against adversarial attacks, which helps promoting trustworthy deep learning applications.
-No potential misuse or negative social impact can be think of.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: No potential misuse or negative social impact can be think of.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We cite all the assets we used properly.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We will provide code and pretrained models with proper documentations.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: We have no human subjects in our paper.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: We have no human subjects in our paper.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@
-Criticality Leveraged Adversarial Training (CLAT) for Boosted Performance via Parameter Efficiency
+Boosting Adversarial Robustness with CLAT: Criticality-Leveraged Adversarial Training
 Abstract
-Adversarial training enhances neural network robustness but suffers from a tendency to overfit and increased generalization errors on clean data.
-This work introduces CLAT, an innovative approach that mitigates adversarial overfitting by introducing parameter efficiency into the adversarial training process, improving both clean accuracy and adversarial robustness.
-Instead of tuning the entire model, CLAT identifies and fine-tunes robustness-critical layers—those predominantly learning non-robust features—while freezing the remaining model to enhance robustness.
-It employs dynamic critical layer selection to adapt to changes in layer criticality throughout the fine-tuning process.
-Empirically, CLAT can be applied on top of existing adversarial training methods, significantly reduces the number of trainable parameters by approximately 95%, and achieves more than a 2% improvement in adversarial robustness compared to baseline methods.
+Adversarial training (AT) is a common technique for enhancing neural network robustness.
+Typically, AT updates all trainable parameters, but such comprehensive adjustments can lead to overfitting and increased generalization errors on clean data.
+Research suggests that fine-tuning specific parameters may be more effective; however, methods for identifying these essential parameters and establishing effective optimization objectives remain unclear and inadequately addressed.
+We present CLAT, an innovative adversarial fine-tuning algorithm that mitigates adversarial overfitting by integrating "criticality" into the training process.
+Instead of tuning the entire model, CLAT identifies and fine-tunes fewer parameters in robustness-critical layers—those predominantly learning non-robust features—while keeping the rest of the model fixed.
+Additionally, CLAT employs a dynamic layer selection process that adapts to changes in layer criticality during training.
+Empirical results demonstrate that CLAT can be seamlessly integrated with existing adversarial training methods, enhancing clean accuracy and adversarial robustness by over 2% compared to baseline approaches.
 1 Introduction
 Advancements in deep learning models have markedly improved image classification accuracy.
-Despite this, their vulnerability to adversarial attacks — subtle modifications to input images that mislead the model — remains a significant concern [9, 28].
-The research community has been rigorously exploring theories to comprehend the mechanics behind adversarial attacks [4].
+Despite this, their vulnerability to adversarial attacks — subtle modifications to input images that mislead the model — remains a significant concern (Goodfellow et al., 2015; Szegedy et al., 2014).
+The research community has been rigorously exploring theories to comprehend the mechanics behind adversarial attacks (Bai et al., 2021).
 Ilyas et al.
-[14] uncover the coexistence of robust and non-robust features in standard datasets.
-Adversarial vulnerability largely stems from the presence of non-robust features in models trained on standard datasets, which, while highly predictive and beneficial for clean accuracy, are susceptible to noise [28].
-Unfortunately, it is observed that deep learning models tend to preferentially learn these non-robust layers.
+(2019) uncover the coexistence of robust and non-robust features in standard datasets.
+Adversarial vulnerability largely stems from the presence of non-robust features in models trained on standard datasets, which, while highly predictive and beneficial for clean accuracy, are susceptible to noise (Szegedy et al., 2014).
+Unfortunately, it is observed that deep learning models tend to preferentially learn these non-robust features.
 Inkawhich et al.
-[15, 16] further demonstrate that adversarial images derived from the hidden features of certain intermediate non-robust/“critical” layers exhibit enhanced transferability to unseen models.
-This suggests a commonality in the non-robust features captured by these layers.
-While identifying these critical layers to improve their robustness is appealing, this process often requires the costly generation of attacks against each individual network layer.
-Efficient methods to identify and effectively address the criticality of such layers are still lacking.
-In contrast to layer-wise feature vulnerability analysis, adversarial training [3, 19, 8], involves training entire neural networks with adversarial examples generated in real-time.
-This approach inherently encourages all layers in the model to learn robust features from adversarial images, thereby enhancing the model’s resilience against attacks.
-However, given the more challenging optimization process of learning from adversarial examples than from clean examples, adversarial training also brings hurdles such as heightened errors on clean data and susceptibility to overfitting, ultimately diminishing its effectiveness in practical applications [25, 36, 23, 17].
-Despite various efforts to enhance adversarial training, such as modifying input data and adjusting loss functions [12, 23, 36, 30, 32, 21], they still frequently fall short in alleviating the aforementioned challenges.
+(2019; 2020) further demonstrate that adversarial images derived from the hidden features of
 Figure 1: CLAT overview.
 CLAT fine-tunes the selected critical layers (red) while freezing other layers (grey).
-fine-tuning objective is computed per Equation 9.
+fine-tuning objective is computed per Eq.
+6.
 Critical layers are adjusted periodically.
 Pseudocode is provided in Appendix A.
+certain intermediate non-robust/“critical” layers exhibit enhanced transferability to unseen models.
+This suggests a commonality in the non-robust features captured by these layers.
+While identifying these critical layers to improve their robustness is appealing, this process often requires the time-consuming generation of attacks against each individual layer.
+Methods to identify and effectively address the criticality of such layers are still lacking.
+In contrast to layer-wise feature vulnerability analysis, adversarial training (Athalye et al., 2018; Madry et al., 2019; Croce & Hein, 2020), involves training entire neural networks with adversarial examples generated in real-time.
+This approach inherently encourages all layers in the model to learn robust features from adversarial images, thereby enhancing the model’s resilience against attacks.
+However, given the more challenging optimization process of learning from adversarial examples than from clean ones, adversarial training also brings hurdles such as heightened errors on clean data and susceptibility to overfitting, ultimately diminishing its effectiveness in practical applications (Schmidt et al., 2018; Zhang et al., 2019; Raghunathan et al., 2019; Javanmard et al., 2020).
+Despite various efforts to enhance adversarial training, such as modifying input data and adjusting loss functions (Hitaj et al., 2021; Raghunathan et al., 2019; Zhang et al., 2019; Wang et al., 2020; Wu et al., 2020b; Pang et al., 2022), these approaches still frequently fall short in alleviating the aforementioned issues.
 In light of these challenges, we introduce CLAT, a paradigm shift in adversarial training, where we mitigate overfitting during adversarial training by identifying and tuning only the robustness-critical model layers.
@@
 Subsequently, our algorithm meticulously fine-tunes these critical layers to remove their non-robust features and reduce their criticality, while freezing the other, non-critical layers.
-Dynamic selection of critical layers is conducted during the training process to always focus the fine-tuning on the most-in-need layers, avoiding the overfitting of full-model adversarial training.
-CLAT therefore enhances both clean and adversarial accuracy compared to previous adversarial training methods.
-Besides reducing overfitting, the critical layer selection in training also makes CLAT a streamlined, parameter-efficient fine-tuning method, enabling memory savings with its significantly reduced trainable parameters.
+Dynamic selection of critical layers is conducted during the training process to always focus fine-tuning on the most-in-need layers, avoiding the overfitting of full-model adversarial training.
+CLAT therefore achieves both clean and adversarial state-of-the-art (SOTA) accuracy compared to previous adversarial training methods.
 In summary, we make the following contributions in this paper:
-  • We introduce the “criticality index”, a quantitative metric designed to efficiently identify critical layers for the adversarial vulnerability of the model.
+  • We introduce the “criticality index”, a quantitative metric designed to identify critical layers for the adversarial vulnerability of a model with minimal overhead.
   • We develop a specialized adversarial training objective focused on reducing the criticality of the identified critical layers to bolster overall model robustness.
-  • We propose CLAT, a parameter-efficient adversarial fine-tuning algorithm focusing on layer criticality reduction.
-CLAT mitigates overfitting in the adversarial training process, and applies to various model training scenarios and baseline adversarial training methods.
-CLAT stands out by markedly reducing overfitting risks, maintaining or even slightly improving the clean accuracy while significantly enhancing adversarial resilience by at least 2% with a 96% reduction in trainable parameters.
+  • We propose CLAT, an adversarial fine-tuning algorithm that mitigates overfitting by focusing on reducing the criticality of fewer than 4% of trainable parameters.
+CLAT integrates seamlessly into diverse model training scenarios and baseline adversarial training methods.
+CLAT stands out by markedly reducing overfitting risks, boosting both clean accuracy and adversarial resilience by approximately 2%.
 2 Related Work
 2.1 Adversarial training
-Adversarial Training (AT) was first introduced by Goodfellow et al.
-[9], who demonstrated how the integration of adversarial examples into the training process could substantially improve model robustness.
-This idea evolved into a sophisticated minimax optimization approach with Projected Gradient Descent Adversarial Training (PGD-AT) [19], which employs PGD attacks in training.
-Regarded as the gold standard in AT, PGD-AT generates adversarial training samples using multiple steps of projected gradient descent, leading to substantially improved empirical robustness [6, 3, 8].
-Further refining this approach, TRADES [36] optimizes a novel loss function to balance classification accuracy with adversarial robustness.
-Recent enhancements in AT, including model ensembling and data augmentation, have also produced notable improvements in model resilience.
-[33, 34, 7]Inkawhich et al.
-[15] propose “Activation Attacks” (AA) which underscore the efficacy of leveraging intermediate model layers for generating stronger adversarial attacks, suggesting that incorporating AA in adversarial training could fortify defenses.
+Adversarial Training (AT) was first introduced by (Goodfellow et al., 2015), who demonstrated how the integration of adversarial examples into the training process could substantially improve model robustness.
+This idea evolved into a sophisticated minimax optimization approach with Projected Gradient Descent Adversarial Training (PGD-AT) (Madry et al., 2019), which employs PGD attacks in training.
+Regarded as the gold standard in AT, PGD-AT generates adversarial training samples using multiple steps of projected gradient descent, leading to substantially improved empirical robustness (Carlini & Wagner, 2017; Athalye et al., 2018; Croce & Hein, 2020).
+Further refining this approach, TRADES (Zhang et al., 2019) optimizes a novel loss function to balance classification accuracy with adversarial robustness.
+Recent enhancements in AT, including model ensemble and data augmentation, have also produced notable improvements in model resilience.
+(Xie et al., 2020; Yang et al., 2020; Carmon et al., 2022)Inkawhich et al.
+(2019) propose “Activation Attacks” (AA) which underscore the efficacy of leveraging intermediate model layers for generating stronger adversarial attacks, suggesting that incorporating AA in adversarial training could fortify defenses.
 Their findings provide a foundation for our method which integrates these intermediate critical layers into our adversarial training strategy.
-2.2 Efficient adversarial training
-Adversarial training methods like PGD-AT and TRADES are computationally intensive, necessitating the multi-step generation of adversarial examples, optimization of complex objective functions, and comprehensive model tuning [26].
-Numerous efforts to enhance the efficiency of adversarial training have been proposed.
-Shafahi et al.
-[26] introduced an efficient model known as “Free” AT, using a single backpropagation step for both training and PGD adversary generation to accelerate adversarial training.
-However, this approach struggled with maintaining model robustness and increased overfitting, primarily due to gradient alignment issues.
-Wong et al.
-[31] then proposed “Fast” AT, which aimed to address these shortcomings, but was later found to be vulnerable to similar problems [2].
-As a result, Andriushchenko and Flammarion [2] developed GradAlign but this approach tripled the training time due to second-order gradient computation.
-Lastly, RiFT leverages network layer redundancies to improve general performance, but its heuristic-based layer redundancy measurement limits final robustness [37].
-In contrast, CLAT uses a theoretically grounded, dynamic, critical layer selection mechanism, resulting in a more robust and parameter-efficient adversarial training algorithm.
-Furthermore, CLAT does not require a fully trained model and can be combined with existing fast-AT methods to enhance performance and mitigate overfitting.
+2.2 Adversarial training improvements and robust overfitting
+Adversarial training methods like PGD-AT and TRADES are computationally expensive and prone to overfitting, requiring multi-step adversary generation, complex objectives, and extensive model tuning (Shafahi et al., 2019).
+To improve efficiency, Shafahi et al.
+(2019) proposed “Free” AT, which accelerates training by using a single backpropagation step for both training and PGD adversary generation.
+However, gradient alignment issues led to reduced robustness and increased overfitting.
+Similarly, Wong et al.
+(2020) introduced “Fast” AT, but it also suffered from similar weaknesses (Andriushchenko & Flammarion, 2020), prompting the development of GradAlign.
+Unfortunately, GradAlign tripled training time due to second-order gradient computation.
+Later efforts sought to address robust overfitting.
+RiFT (Zhu et al., 2023) improved general performance by leveraging layer redundancies but was constrained by heuristic redundancy measurements.
+Zhang et al.
+(2024) mitigated overfitting by disentangling natural and adversarial objectives, yet model-wide adjustments still limited robustness.
+In contrast, CLAT uses a theoretically grounded, dynamic, critical layer selection mechanism, resulting in improved robust generalization by tuning a critical subset of layers.
+Furthermore, CLAT is agnostic to attack generation methods in the AT processing, making it an ideal complement to existing fast-AT methods to mitigate overfitting.
 3 Methods
-Building on prior attack and defense research [15, 16, 37] which demonstrates that not all model layers equally learn non-robust features, we aim to improve model robustness and parameter efficiency by identifying and fine-tuning only those critical layers that are prone to learning non-robust features, while keeping the non-critical layers frozen.
+Building on prior attack and defense research (Inkawhich et al., 2019; 2020; Zhu et al., 2023) which demonstrates that not all model layers equally learn non-robust features and having all layers learn robust features leads to overfitting, we aim to improve model robustness by identifying and fine-tuning only those critical layers that are prone to learning non-robust features, while keeping the non-critical layers frozen.
 In this section, we begin by defining and identifying critical layers, then outline our objectives for reducing their criticality.
-Finally, we present our complete CLAT algorithm, which effectively enhances model robustness.
+Finally, we present our complete CLAT algorithm, which effectively mitigates overfitting and enhances model robustness.
 3.1 Layer Criticality
@@
 Under adversarial perturbation, features from all layers will be altered, leading to incorrect outputs.
-To understand how the learned parameters of the model affect the features under perturbation, we consider a function $G_{i}(\cdot)$ defined at the output of layer $i$ on inputs close to a clean data point $x$:
-$G_{i}(z)=||F_{i}(z)-F_{i}(x)||_{2}^{2}.$ (2)
-Clearly, $z=x$ is a minima of $G_{i}(z)$. The worst-case curvature of the function $G_{i}$ at the neighborhood of $z=x$ can be estimated following the formulation by Moosavi-Dezfooli et al. [20] as
-$\nu_{i}(x)=\frac{\nabla G_{i}(x^{\prime})-\nabla G_{i}(x)}{||x^{\prime}-x||_{2}}=\frac{\nabla G_{i}(x^{\prime})}{||x^{\prime}-x||_{2}},$ (3)
-where $x^{\prime}$ is a worst-case perturbation (adversarial attack) maximizing $G_{i}(z)$ in the vicinity of $x$, and $\nabla G_{i}(x)=0$ by definition given it is a minimum.
-The numerator in the curvature formulation can be further derived as
-$\nu_{i}(x)=\frac{\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}^{T}(F_{i}(x^{\prime})-F_{i}(x))}{||x^{\prime}-x||_{2}}.$ (4)
-Following the observation in [20], a higher curvature indicates the feature to be more sensitive to adversarial examples.
-We can therefore use the curvature formulation $\nu_{i}(x)$ under a fixed perturbation budget $||x^{\prime}-x||_{p}\leq\epsilon$ to estimate the layer sensitivity, or weakness.
-In practice, it is difficult to explicitly instantiate $\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}$ for a neural network.
-To this end, we simplify the formulation in Equation 4 by assuming $\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}$ as a uniform vector.
+Following previous work (Hein & Andriushchenko, 2017; Finlay et al., 2018), the robustness, or weakness, of the feature can be linked to the local Lipschitz constant of function $F_{i}(\cdot)$.
+For a easier computation, we consider the worst-case feature difference under a fixed input perturbation budget of $\epsilon$.
 This leads to our definition of the $\epsilon$-weakness of layer $i$’s feature as:
@@
 where $N_{i}$ denotes the dimensionality of the output features at layer $i$, therefore normalizing the weakness measurement of layers with different output sizes.
-The weakness measurement is proportional to the curvature estimation in Equation 4.
+The weakness measurement is proportional to the local Lipschitz constant.
 A higher weakness value indicates that the feature vector is more vulnerable to input perturbations.
 The functionality of cascading layers from 1 to $i$ affects the vulnerability of the hidden features, as described by this formulation.
-For the purpose of efficient fine-tuning, we want to identify the layers that are the most critical to the lack of robustness, characterized by their increased susceptibility to perturbations from adversarial inputs.
+Alternatively, Moosavi-Dezfooli et al.
+(2019) suggests the local curvature can be a more accurate estimation of robustness than the Lipschitz constant.
+However, the computation and optimization of curvature involves costly higher-order gradient computation.
+We provide additional derivations in Appendix E to show the feature weakness defined in Eq.
+2 is also an effective approximation to the local curvature value.
+For the purpose of mitigating overfitting, we want to identify the layers that are the most critical to the lack of robustness, characterized by their increased susceptibility to perturbations from adversarial inputs.
+Other already-robust layers shall then be fixed to avoid further overfitting to the adversarial training objective.
 We therefore provide the following definition:
 Definition 3.1.
-Critical layer: A layer is considered critical if it exhibits a greater propensity to learn non-robust features or demonstrates heightened sensitivity to adversarial input perturbations relative to other layers in the model.
+Critical layer: A layer is considered critical if it exhibits a greater propensity to learn non-robust features or demonstrates diminished robustness to adversarial input perturbations relative to other layers in the model.
 To this end, we single out the contribution of each layer’s functionality to the weakness of the features after it with a Layer Criticality Index $\mathcal{C}_{f_{i}}$, which is formulated as
@@
 As a sanity check, the feature weakness at the output of layer $i$ can be attributed to the criticality of all previous layers as $\mathcal{W}_{\epsilon}(F_{i})=\prod_{k=1}^{i}\mathcal{C}_{f_{k}}$.
-Conversely, a layer with a larger criticality index will increase the weakness of the features after it, indicating the layer is critical according to Definition 3.1, as it heightens the feature sensitivity to the adversarial input.
-One drawback of the formulation in Equation (6) is that computing the feature weakness involves finding the worst-case perturbation against the hidden features at each layer, which is a costly process to conduct sequentially for all layers.
-In practice, we find it possible to approximate the worst-case perturbation against features with an untargeted PGD attack against the model output, so that we can use the same PGD perturbation $\delta$ to estimate the feature weakness of all layers following Equation 5.
+Conversely, a layer with a larger criticality index will increase the weakness of the features after it, indicating the layer is critical according to Definition 3.1, as it weakens the feature robustness to the adversarial input.
+One drawback of the formulation in Eq.
+3 is that computing the feature weakness involves finding the worst-case perturbation against the hidden features at each layer, which is a costly process to conduct sequentially for all layers.
+In practice, we approximate the worst-case perturbation against features with an untargeted PGD attack against the model output, so that we can use the same PGD perturbation $\delta$ to estimate the feature weakness of all layers following Eq.
+2.
 In this way, with a reasonably sufficient batch size, we can compute the critical indices for all layers in a model with two forward passes: one with the clean input $x$ and one with the PGD attack input $x+\delta$.
@@
 For a critical layer $i$, we optimize the trainable parameters to minimize $\mathcal{C}_{f_{i}}$.
-Note that in the criticality formulation in Equation 6, the weakness of the previous layer’s output, $\mathcal{W}_{\epsilon}(F_{i-1})$, is constant with respect to $f_{i}$.
+Note that in the criticality formulation in Eq.
+3, the weakness of the previous layer’s output, $\mathcal{W}_{\epsilon}(F_{i-1})$, is constant with respect to $f_{i}$.
 Thus, the optimization objective for $f_{i}$ can be simplified as
@@
 where a single perturbation is utilized to capture the weakness across all critical layers.
-In practice, a projected gradient ascent optimization with random start is used to solve the inner maximization.
-Minimizing the objective in Equation 8 by adjusting the trainable variables of the critical layers will reduce their feature weaknesses.
+In the common setting, a projected gradient ascent optimization with random start is used for the inner maximization.
+Minimizing the objective in Eq.
+5 by adjusting the trainable variables of the critical layers will reduce their feature weaknesses.
 However, the removal of non-robust features in these layers may affect the functionality of the model on clean inputs.
@@
 where the hyperparameter $\lambda$ serves as a balancing factor between the two loss terms.
-Note that only the selected critical layers $f_{S}$ are optimized in Equation 9 while the other non-critical layers are frozen, justifying our fine-tuning to be parameter-efficient.
+Note that only the selected critical layers $f_{S}$ are optimized in Eq.
+6 while the other non-critical layers are frozen, preventing them from further overfitting.
 3.3 CLAT adversarial training
-Similar to parameter-efficient fine-tuning techniques in other domains, we design CLAT as a fine-tuning approach, which is applied to neural networks that have undergone preliminary training.
+We design CLAT as a fine-tuning approach, which is applied to neural networks that have undergone preliminary training.
 The pretraining phase allows all layers in the model to capture useful features, which will facilitate the identification of critical layers in the model.
@@
 We then fine-tune critical layers only while freezing the rest of the layers.
-This process is illustrated in Figure 1, and the pseudocode is provided in Algorithm 1 in Appendix A for greater clarity.
-Compared to full-model fine-tuning, CLAT not only improves parameter efficiency with fewer trainable parameters, but also reduces overfitting, a common issue in standard adversarial training.
+This process is illustrated in Fig.
+1, and the pseudocode is provided in Algorithm 1 in Appendix A for greater clarity.
 As fine-tuning progresses, the critical layers will be updated to reduce their criticality, making them less critical than some of the previously frozen layers.
@@
 We conducted experiments using two widely recognized image classification datasets, CIFAR10 and CIFAR100.
-Each dataset includes 60,000 color images, each 32×32 pixels, divided into 10 and 100 classes respectively [18].
-For our experiments, we deployed a suite of network architectures: WideResnets (34-10, 70-16) [35], ResNets (50, 18) [11], DenseNet-121 [13], PreAct ResNet-18[10], and VGG-19 [27].
+Each dataset includes 60,000 color images, each 32×32 pixels, divided into 10 and 100 classes respectively (Krizhevsky & Hinton, 2009).
+For our experiments, we deployed a suite of network architectures: WideResnets (34-10, 70-16) (Zagoruyko & Komodakis, 2017), ResNets (50, 18) (He et al., 2016b), DenseNet-121 (Huang et al., 2017), PreAct ResNet-18(He et al., 2016a), and VGG-19 (Simonyan & Zisserman, 2015).
 In this paper, these architectures are referred to as WRN34-10, WRN70-16, RN50, RN18, DN121, Preact RN18 and VGG19 respectively.
 Training and evaluation
-Since CLAT can be layered over clean pretraining, partial training, or other adversarial methods, results incorporating CLAT are denoted in our tables as "X+ CLAT," where “X” specifies the baseline method used prior to applying CLAT.
-Typically, this baseline method was applied for the first 50 epochs, followed by a fine-tuning phase during which CLAT was run for an additional 50 epochs.
-Models trained exclusively with PGD-AT typically require 150 epochs to achieve reported performances.
-We use PGD for attack generation unless stated otherwise in the baseline method, which employs a random start [19], with an attack budget of $\epsilon$ = 0.03 under the $\ell_{\infty}$ norm, an update step size of $\alpha$ = 0.007, and 10 attack steps.
+Since CLAT can be layered over clean pretraining, partial training, or other adversarial methods, results incorporating CLAT are denoted in our tables as "X + CLAT," where "X" refers to the baseline method applied prior to CLAT.
+Typically, this baseline method is run for the first 50 epochs, followed by fine-tuning during which CLAT is applied for an additional 50 epochs.
+The total number of epochs is in line with the 100 epochs used in previous PGD-based adversarial training work (Zhang et al., 2019; Zhu et al., 2023).
+For our baseline, we use PGD for attack generation during training, following a random start (Madry et al., 2019), with an attack budget of $\epsilon=0.03$ under the $\ell_{\infty}$ norm, a step size of $\alpha=0.007$, and 10 attack steps.
+The same settings apply to PGD attack evaluations.
+AutoAttack evaluations (Croce & Hein, 2020) also use a budget of $\epsilon=0.03$ under the $\ell_{\infty}$ norm, with no restarts for untargeted APGD-CE, 9 target classes for APGD-DLR, 3 target classes for Targeted FAB, and 5000 queries for Square Attack.
+These settings remain consistent unless explicitly noted otherwise.
 Experiments were conducted on a Titan XP GPU, starting with an initial learning rate of 0.1, which was adjusted according to a cosine decay schedule.
@@
 White-box robustness
-Table 1 presents the clean accuracy and white-box $\ell_{\infty}$ PGD [6] accuracy across various adversarial training techniques.
-The versatility and effectiveness of CLAT are demonstrated by combining it with various standard adversarial training methods.
-CLAT mitigates the overfitting commonly observed in traditional adversarial training methods, improving both clean and adversarial accuracy compared to the corresponding baselines.
-We also show that reducing trainable parameters alone does not guarantee a performance improvement by comparing CLAT against other parameter-efficient fine-tuning methods like LoRA [1] and RiFT [37].
-Our theoretically-grounded criticality indices enable the precise identification of critical layers and the focused elimination of non-robust features from these pivotal layers.
-Furthermore, we verify that fast adversarial training techniques as in [31] can be applied to solve the inner maximization in the CLAT training objective in Equation 8, where the “CLAT (Fast)” method also improves performance and robustness over Fast-AT baselines.
-Besides the main table, we provide additional robustness results in Appendix B, where we showcase CLAT’s improved robustness against a variant of the white-box Auto Attack [8] in LABEL:wb_aa and verify the generalizability of the robustness to attacks of various strengths in Figure 4.
+Table 1 and Table 2 present white-box evaluation results using the PGD and Auto attack frameworks, respectively.
+These tables illustrate CLAT’s versatility and effectiveness when combined with various standard adversarial training methods, including state-of-the-art benchmarks from RobustBench (Croce et al., 2020).
+CLAT effectively mitigates the overfitting typically observed in traditional adversarial training, thereby enhancing both clean and adversarial accuracy compared to baseline methods.
+Table 1 further highlights that reducing trainable parameters alone does not necessarily lead to improved performance.
+CLAT surpasses other parameter-efficient fine-tuning methods, such as LoRA (Aleem et al., 2024) and RiFT (Zhu et al., 2023), due to its ability to precisely identify critical layers and eliminate non-robust features from these layers.
+Additionally, we show that fast adversarial training techniques, as discussed by Wong et al.
+(2020), can be applied to address the inner maximization problem in the CLAT training objective described in Eq.
+5.
+The "CLAT (Fast)" method not only enhances performance but also improves robustness compared to Fast-AT baselines.
+Notably, CLAT models are trained with PGD-like attacks on hidden features without seeing Auto Attacks directly, but their robustness persists under these attacks (see Table 2).
+This suggests that different attacks across various networks share similarities in exploiting non-robust features.
+By addressing these non-robust features through critical layer fine-tuning, CLAT ensures that the robustness is adaptable across different attack settings and models.
+Lastly, we assess the robustness across various attack strengths in Appendix B.
 Table 1: Comparative performance of CLAT across various networks and adversarial training/fine-tuning techniques. Robustness is evaluated with white-box PGD attacks.
@@
 clean adversarial clean adversarial
-DN121 pgd-at [19] 80.05 58.15 57.18 31.76
+DN121 pgd-at (Madry et al., 2019) 80.05 58.15 57.18 31.76
 pgd-at + clat 81.03 60.60 58.79 33.23
-WRN70-16 Peng et al. [22] 93.27 71.07 70.20 42.61
+WRN70-16 (Peng et al., 2023) 93.27 71.07 70.20 42.61
 peng et al. + clat 93.56 72.25 71.94 44.12
-Bai et al. [5] 92.23 64.55 69.17 40.86
+(Bai et al., 2024) 92.23 64.55 69.17 40.86
 bai et al. + clat 92.77 64.92 70.17 41.64
@@
 WRN34-10 pgd-at 87.41 55.40 59.19 31.66
-pgd-at + lora [1] 73.36 56.17 55.56 31.43
-pgd-at + rift [37] 87.89 55.41 62.35 31.64
+pgd-at + lora (Aleem et al., 2024) 73.36 56.17 55.56 31.43
+pgd-at + rift (Zhu et al., 2023) 87.89 55.41 62.35 31.64
 pgd-at + clat 88.97 57.11 62.38 32.05
-trades [36] 87.60 56.61 60.56 31.85
+trades (Zhang et al., 2019) 87.60 56.61 60.56 31.85
 trades + rift 87.55 56.72 61.01 32.03
@@
 trades + clat 81.89 54.57 58.82 31.06
-mart [29] 76.77 56.90 51.46 31.47
+mart (Wang et al., 2019) 76.77 56.90 51.46 31.47
 mart + rift 77.14 56.92 52.42 31.48
 mart + clat 76.82 57.55 53.01 33.23
-awp [32] 78.40 53.83 52.85 31.00
+awp (Wu et al., 2020b) 78.40 53.83 52.85 31.00
 awp + rift 78.79 53.84 54.89 31.05
 awp + clat 79.01 55.27 55.39 32.08
-score [21] 84.20 54.59 54.83 29.49
+score (Pang et al., 2022) 84.20 54.59 54.83 29.49
 score + rift 85.65 54.62 57.63 29.50
 score + clat 86.11 55.78 57.66 30.23
-PREACT RN18 fast-at [31] 81.46 45.55 50.10 27.72
+autolora (Zhang et al., 2024) 84.2 54.27 62.1 32.71
+autolora + clat 86.45 58.53 64.91 35.82
+PREACT RN18 fast-at (Wong et al., 2020) 81.46 45.55 50.10 27.72
 fast-at + clat 84.46 52.13 54.33 29.22
 fast-at + clat (fast) 82.72 49.62 52.10 27.99
+Table 2: Adversarial Accuracy on CIFAR-10 and CIFAR-100 when subjected to AutoAttack (AA).
+Network Method Adversarial Accuracy (%)
+CIFAR-10 CIFAR-100
+DN121 pgd-at 47.56 23.13
+pgd-at + clat 49.91 25.74
+WRN70-16 pgd-at 54.32 28.25
+pgd-at + clat 57.64 30.98
+(Carlini & Wagner, 2017) 66.1 -
+(Carlini & Wagner, 2017) + clat 68.43 -
+RN50 pgd-at 46.22 23.48
+pgd-at + clat 49.45 25.81
+(Madry et al., 2019) 49.25 -
+(Madry et al., 2019) + clat 51.64 -
+WRN34-10 pgd-at 51.50 25.56
+pgd-at + clat 52.88 27.62
+(Croce & Hein, 2020) 57.3 -
+(Croce & Hein, 2020) + clat 59.98 -
+VGG19 pgd-at 40.42 19.54
+pgd-at + clat 41.72 20.45
+RN18 pgd-at 40.48 20.21
+pgd-at + clat 42.86 21.76
+twins (Liu et al., 2023) 47.89 25.45
+twins + clat 51.39 28.12
+autolora (Zhang et al., 2024) 48.95 27.48
+autolora + clat 53.21 30.49
 Black-box robustness
-Table 2 evaluates the robustness against black-box attacks between models trained solely using PGD-AT and those augmented with CLAT.
+Table 3 and Table 4 evaluate the robustness against black-box attacks (Auto Attack and PGD-AT respectively) between models trained solely using PGD-AT and those augmented with CLAT.
 Attack settings are the same as those of the white-box attacks.
@@
 More significantly, models trained with CLAT consistently outperform those trained with PGD-AT, maintaining superior resilience in both black-box and white-box settings, regardless of the attack method or models employed.
+Table 3: Comparative Analysis of Black-box Auto Attack (AA) Accuracy on CIFAR-10 and CIFAR-100. Model in each row is the attacker and each column the victim.
+Network Method CIFAR-10 Adv. Acc. (%) CIFAR-100 Adv. Acc. (%)
+DN121 RN50 VGG19 RN18 DN121 RN50 VGG19 RN18
+DN121 pgd-at - 52.50 44.21 45.45 - 27.64 23.21 24.26
+pgd-at + clat - 55.83 47.53 48.92 - 29.89 26.71 26.93
+RN50 pgd-at 54.23 - 43.56 43.24 27.91 - 23.02 23.51
+pgd-at + clat 56.72 - 46.21 47.01 30.11 - 26.34 25.86
+VGG19 pgd-at 55.32 55.45 - 46.72 27.84 28.15 - 24.20
+pgd-at + clat 59.83 59.72 - 49.31 30.20 29.79 - 26.55
+RN18 pgd-at 53.21 51.73 43.21 - 29.31 26.75 22.91 -
+pgd-at + clat 56.75 54.45 46.53 - 31.25 29.41 25.95 -
 Table 4: Comparative Analysis of Black-box PGD Accuracy on CIFAR-10 and CIFAR100. Model in each row is the attacker and each column the victim.
@@
 Rn50 pgd-at - 74.83 68.01 67.44 - 46.82 40.55 40.10
-clat - 76.45 71.25 70.12 - 48.49 44.34 43.91
+pgd-at + clat - 76.45 71.25 70.12 - 48.49 44.34 43.91
 Dn121 pgd-at 72.24 - 69.53 68.38 44.45 - 40.63 41.22
-clat 74.55 - 71.78 70.56 46.78 - 43.62 42.88
+pgd-at + clat 74.55 - 71.78 70.56 46.78 - 43.62 42.88
 Vgg19 pgd-at 65.72 67.56 - 62.26 47.86 46.56 - 40.55
-clat 66.46 70.72 - 65.78 49.25 48.72 - 42.72
+pgd-at + clat 66.46 70.72 - 65.78 49.25 48.72 - 42.72
 Rn18 pgd-at 74.82 70.21 61.83 - 46.28 45.59 39.21 -
-clat 76.23 73.19 63.96 - 48.89 47.72 41.78 -
+pgd-at + clat 76.23 73.19 63.96 - 48.89 47.72 41.78 -
 4.3 Ablation Studies
@@
 Here, we analyze how the number of pretraining epochs affects CLAT performance.
-Figure 2 shows the training curves for different allocations of PGD pretraining epochs and CLAT fine-tuning epochs within a 100-epoch training budget.
-The overfitting of PGD-AT is evident as adversarial accuracy plateaus and declines towards the end, as documented in previous research [24].
+Fig.
+2 shows the training curves for different allocations of PGD pretraining epochs and CLAT fine-tuning epochs within a 100-epoch training budget.
+The overfitting of PGD-AT is evident as adversarial accuracy plateaus and declines towards the end, as documented in previous research (Rice et al., 2020).
 In contrast, CLAT continues to improve adversarial accuracy, effectively addressing this issue.
@@
 This phenomenon supports our theoretical insight that criticality can be linked to the curvature of the local minima to which each layer converges during adversarial training.
-Flattening the local minima of critical layers aids in model generalization.
-We plan to explore this phenomenon further in future work.
-Figure 2: Comparative analysis of CLAT performance on WRN34-10: Clean and adversarial accuracy on CIFAR-10 across partially trained models. Note: PGD-AT requires 150 epochs to fully converge.
+Figure 2: Comparative analysis of CLAT performance on WRN34-10: Clean and adversarial accuracy on CIFAR-10 across partially trained models.
+Table 5: Adversarial accuracy under PGD attack: Comparison of PGD-AT, PGD-AT + CLAT, and PGD-AT + CLAT (Non-dynamic) on CIFAR-10 and CIFAR-100.
+Method CIFAR-10 CIFAR-100
+DN121 RN50 DN121 RN50
+pgd-at 58.15 56.35 31.76 33.01
+clat 60.60 59.54 33.23 36.23
+clat (nd) 57.01 54.22 30.34 32.98
+Table 6: Adversarial accuracy under Auto attack: Comparison of PGD-AT, PGD-AT + CLAT, and PGD-AT + CLAT (Non-dynamic) on CIFAR-10 and CIFAR-100.
+Method CIFAR-10 CIFAR-100
+DN121 RN50 DN121 RN50
+pgd-at 47.56 46.22 23.13 23.48
+clat 49.91 49.45 25.74 25.81
+clat (nd) 47.12 46.08 22.26 22.91
+4.3.2 Ablating on critical layer selection
+The choice of critical layer selection is another important feature impacting the performance of CLAT.
+We begin by examining the effect of dynamic layer selection.
+Table 6 and Table 6 highlight that dynamic selection is crucial to CLAT’s performance.
+Using the same layers throughout the process tends to cause overfitting and results in lower accuracies compared to the PGD-AT baseline.
+To verify the significance of the selected critical layers, we compare CLAT with an alternative approach in which random layers are dynamically selected for fine-tuning instead of the critical layers.
+The results of this comparison are detailed in Table 7 and Table 12 (in Appendix C).
+The data demonstrates that selecting critical layers significantly enhances the model’s adversarial robustness and clean accuracy.
+This observation is bolstered by our ablation study in Appendix Table 20 illustrating the performance effect of choosing the smallest versus largest critical indices for fine-tuning.
+Furthermore, Table 9 indicates near-identical critical layer selections within the same model, even across diverse datasets.
+This evidence supports our assertion that the variation in layer criticality arises from inherent properties within the model architecture, where certain layers are predisposed to learning non-robust features.
+Table 7: Ablating CLAT Layer choices on CIFAR-10: The columns present PGD-10 and Auto Attack (AA) evaluation adversarial accuracies for models trained with CLAT, where layers are selected based on criticality or randomly.
+Network Critical Layers Random Layers
+Clean Acc. PGD-10 AA Clean Acc. PGD-10 AA
+DN121 81.03 60.60 49.91 78.85 51.35 39.81
+RN50 83.78 59.54 49.45 79.01 51.44 40.29
+RN18 83.89 55.37 42.86 78.02 51.03 33.50
 Table 8: Top-5 criticality indices by model and dataset. Layers used in CLAT are bolded.
@@
 RN18 11, 10, 4, 2, 12 12, 9, 5, 2, 13
-Table 4: Comparison of PGD-AT, PGD-AT + CLAT, and PGD-AT + CLAT (Non-dynamic) on CIFAR-10 Adv. Accuracy.
-Method DN121 RN50
-pgd-at 58.15 56.35
-clat 60.60 59.54
-clat (nd) 57.01 54.22
-4.3.2 Ablating on critical layer selection
-The choice of critical layer selection is another important feature impacting the performance of CLAT.
-We begin by examining the effect of dynamic layer selection.
-Table 4 highlights that dynamic selection is crucial to CLAT’s performance.
-Using the same layers throughout the process tends to cause overfitting and results in lower accuracies compared to the PGD-AT baseline.
-To verify the significance of the selected critical layers, we compare CLAT with an alternative approach in which random layers are dynamically selected for fine-tuning instead of the critical layers.
-The results of this comparison are detailed in Table 6 and Table 10 (in Appendix C) on CIFAR-10 and CIFAR-100, respectively.
-The data demonstrates that selecting critical layers significantly enhances the model’s adversarial robustness and yields a marginal increase in clean accuracy.
-Moreover, Table 4 indicates near-identical critical layer selections within the same model even across diverse datasets.
-This evidence supports our assertion that the variation in layer criticality stems from inherent properties within the model architecture, wherein certain layers are predisposed to learning non-robust features.
-Table 5: Critical vs. random layers for CLAT on CIFAR-10.
-Network CIFAR-10
-Critical Layers Random Layers
-Clean Acc. Adv. Acc. Clean Acc. Adv. Acc.
-DN121 81.03 60.60 78.85 51.35
-RN50 83.78 59.54 79.01 51.44
-RN18 83.89 55.37 78.02 51.03
 Table 9: Trainable Parameters during CLAT in Various Networks
@@
 rn18 11.2M 590K 5.2%
-vgg19 39.3M 236K 6.0%
-Lastly, we ablate on the number of critical layers used in CLAT for fine-tuning.
-Figure 3 illustrates the trade-off between adversarial test accuracy and the quantity of critical layers selected for fine-tuning.
-Allowing more layers to be fine-tuned enhances the model flexibility, which initially improves CLAT performance.
-However, fine-tuning an excessive proportion of critical layers diminishes CLAT’s effectiveness.
-This is likely due to the diversion of attention towards fine-tuning less-critical layers, detracting from more critical ones.
-This pattern highlights the crucial influence of specific layers on network robustness and suggests a need for deeper research to understand the roles and dynamics of individual layers influencing network robustness.
-Figure 3: Comparative analysis on CLAT performance/adversarial accuracy with respect to number of critical layers used during CLAT
-4.4 Efficiency and stability analysis
-To verify the significance of the selected critical layers, we contrast CLAT with an alternative approach in which random layers are dynamically selected for fine-tuning instead of the critical layers.
-The results of this comparison are detailed in Table 6 and Table 10 (in Appendix C) on CIFAR-10 and CIFAR-100, respectively.
-The cost of optimizing the CLAT training objective in Equation 9 is similar to that of the standard adversarial training given its min-max formulation.
+vgg19 39.3M 236K 0.6%
+Lastly, we conduct an ablation study on the number of layers used in CLAT for fine-tuning.
+Fig.
+3 and Fig.
+7 illustrate the trade-offs between adversarial accuracy and the number of layers selected, as well as clean accuracy and the number of layers, respectively.
+Interestingly, both adversarial and clean accuracy are optimized with the same number of layers.
+Initially, fine-tuning more layers enhances model performance by increasing flexibility; however, this eventually diminishes CLAT’s effectiveness, likely because attention is diverted to less crucial layers at the expense of more important ones.
+This pattern underscores the critical role of specific layers in network robustness and emphasizes the need for deeper research into the dynamics of individual layers.
+Furthermore, we highlight that although the number of layers chosen impacts the learned robustness, CLAT still achieves robustness gains over the early-stopping baseline (no fine-tuning) with up to 10% of the layers selected, demonstrating its stability under small variations in the number of selected layers.
+Figure 3: Comparative analysis on CLAT performance/PGD-10 adversarial accuracy with respect to number of critical layers used during CLAT
+4.4 Overhead and stability analysis
+The cost of optimizing the CLAT training objective in Eq.
+6 is similar to that of the standard adversarial training given its min-max formulation.
 Here, we show that the computational overhead for determining critical indices is negligible.
@@
 We further conducted over 1000 runs for each network to randomly select the data used for criticality estimation, where we find remarkable consistency in computed critical layers, differing in less than 5% of cases, typically involving only one layer change among the top five critical layers.
-The stability means we can use a mere 0.0002% of the training data for criticality estimation every 10 epochs, which introduces neglectable additional complexity to the training process as indicated by the time measurement in the table.
+The stability means we can use a mere 0.0002% of the training data for criticality estimation every 10 epochs, which introduces a neglectable 0.4% additional time to the standard adversarial training process as measured in Table 10.
 Table 10: DenseNet-121 critical layers identified with different amount of data. Time taken to compute critical layers evaluated on TITAN RTX GPU. As a reference, 1 PGD-AT epoch takes 67s.
@@
 5 Conclusions
-In this work, we introduce CLAT, an innovative adversarial training approach that addresses adversarial overfitting issues by fine-tuning only the critical layers vulnerable to adversarial perturbations.
+In this work, we introduce CLAT, an innovative adversarial training approach that addresses robust overfitting issues by fine-tuning only the critical layers vulnerable to adversarial perturbations.
 This method not only emphasizes layer-specific interventions for enhanced network robustness but also sheds light on the commonality in non-robust features captured by these layers, offering a targeted and effective defense strategy.
-Our results reveal that CLAT reduces trainable parameters by about 95% while ensuring significant improvements in clean accuracy and adversarial robustness across diverse network architectures and baseline adversarial training methods.
+Our results reveal that CLAT effectively selects less than 4% critical trainable variables to achieve significant improvements in clean accuracy and adversarial robustness across diverse network architectures and baseline adversarial training methods.
 We limit the scope of this work to improving the empirical robustness of the model by utilizing the critical layers.
@@
 6:         $x+\delta\leftarrow$ PGD attack against $F$
-7:         Compute $\mathcal{W}_{\epsilon}(F_{i})$ for all layers with Equation 5
-8:         Compute $\mathcal{C}_{f_{i}}$ for all layers with Equation 6
+7:         Compute $\mathcal{W}_{\epsilon}(F_{i})$ for all layers with Eq. 2
+8:         Compute $\mathcal{C}_{f_{i}}$ for all layers with Eq. 3
 9:         Critical layers $\mathcal{S}\leftarrow TopK(\mathcal{C}_{f_{i}})$
@@
 12:     for $x,y$ in $minibatches$ do
-13:         Perturbation $\delta\leftarrow$ Equation 8 inner maximization
-14:         $\mathcal{L}_{C}(f_{\mathcal{S}})\leftarrow$ Equation 8
-15:         Weight update $w[\mathcal{S}]$ with Equation 9
+13:         Perturbation $\delta\leftarrow$ Eq. 5 inner maximization
+14:         $\mathcal{L}_{C}(f_{\mathcal{S}})\leftarrow$ Eq. 5
+15:         Weight update $w[\mathcal{S}]$ with Eq. 6
 Appendix B Additional experiment results
-Table LABEL:wb_aa showcases results under a variant of the white-box Auto Attack [8] with an epsilon value of 0.03 and a single restart in Table 1.
-Specifically, we report adversarial accuracies of the “APGDCE” untargeted attack, a component of the Auto Attack ensemble, with similar trends observed for other attacks in the ensemble.
-Although CLAT models employ PGD-like attacks on hidden features during the adversarial training phase and have not seen Auto Attacks, the robustness remains under Auto Attacks.
-This highlights the similarity in non-robust features utilized by different attacks in different networks.
-As CLAT resolves the non-robust features with critical layer fine-tuning, the resulting robustness is generalizable across attack settings and source models of the attacks.
-Table 8: Adversarial accuracy under the white-box “APGDCE” variant of Auto Attack: Models trained using CLAT consistently outperform those trained exclusively with PGD-AT.
-Dataset Network PGD-AT(%) PGD-AT + CLAT (%)
-CIFAR-10 DN121 49.26 50.58
-WRN70-16 54.48 56.90
-RN50 48.54 50.65
-WRN34-10 52.21 54.73
-VGG19 45.35 47.01
-RN18 46.99 48.20
-CIFAR-100 DN121 24.67 26.31
-WRN70-16 30.27 33.32
-RN50 25.44 26.64
-WRN34-10 27.62 31.53
-VGG19 22.50 24.02
-RN18 23.81 24.43
 We further compare the CLAT model robustness with the robustness of the SAT model against white-box attacks of various strengths.
-As illustrated in  Figure 4, though both models are only trained against an attack of one strength ($\epsilon$ = 0.03), the improved robustness of CLAT is consistent across the full spectrum of attack strengths.
+As illustrated in  Fig.
+4, though both models are only trained against an attack of one strength ($\epsilon$ = 0.03), the improved robustness of CLAT is consistent across the full spectrum of attack strengths.
 This shows that CLAT is not overfitting to the specific attack strength used in training.
@@
 Here in Table 12, we provide additional results contrasting CLAT with an alternative approach where random layers are dynamically selected for fine-tuning instead of the critical layers on the CIFAR-100 dataset.
-Table 10: Ablating the layer choices for CLAT fine-tuning on CIFAR-100.
-Network CIFAR-100
-Critical Layers Random Layers
-Clean Acc. Adv. Acc. Clean Acc. Adv. Acc.
-DN121 58.79 33.23 54.85 23.32
-RN50 61.88 36.23 56.50 22.45
-RN18 59.22 32.04 52.43 22.60
+Table 12: Ablating CLAT layer choices on CIFAR-100: The columns present PGD-10 and AutoAttack (AA) evaluation adversarial accuracies for models trained with CLAT, where layers are selected based on criticality or randomly.
+Network Critical Layers Random Layers
+Clean Acc. PGD-10 AA Clean Acc. PGD-10 AA
+DN121 58.79 33.23 25.74 50.45 25.48 20.21
+RN50 61.88 36.23 25.81 52.20 25.51 20.45
+RN18 50.98 28.41 20.45 43.15 20.24 15.89
+Appendix D Rebuttal Experiments
+We thank the constructive feedback from all reviewers.
+We consolidate all additional experimental results suggested by the reviewers in this appendix section.
+We will integrate these results and our discussions in the rebuttal text into the finalized version.
+D.1 Full Training Curves
+Figure 5: White-box PGD-10 adversarial accuracy (y-axis) on CIFAR-10 for WRN34-10 models trained with CLAT fine-tuning starting at Epoch 70 (red), CLAT from scratch (orange), and PGD-AT (blue).
+The learning rate decays to 0 by Epoch 150.
+D.2 Reduced Learning Rate Performance
+Please see Fig. 6.
+Figure 6: White-box accuracies (y-axis) for WRN34-10 on CIFAR-10 for models trained with the original learning rate multiplied by 0.1 at Epoch 70, using CLAT (green) and PGD-AT (red), compared to normal PGD-AT learning rate (purple).
+D.3 Critical Index Variation Over Time
+Please see Table 13.
+Table 13: Critical layers identified at different epochs for various networks.
+Network Ep. 70 Ep. 80 Ep. 90
+DN121 [39, 14, 1, 3, 88] [38, 1, 5, 88, 15] [1, 5, 88, 2, 15]
+RN50 [34, 41, 48] [48, 3, 36] [36, 2, 40]
+RN18 [11] [11] [4]
+D.4 Clean Accuracy and Critical Layer Selection
+Figure 7: Comparative analysis on CLAT performance clean accuracies with respect to the number of critical layers used during CLAT.
+D.5 Ablation Study: Dynamic and Fixed Critical Indices
+Please see Table 14.
+Table 14: Comparison of CIFAR-100 Clean and Adversarial Accuracies (PGD-10) on Different Networks with Fixed and Dynamic Layers for CLAT.
+Model Fixed Layers Dynamic Layers
+DN121 Clean 53.21 58.79
+Adv 39.45 44.12
+RN50 Clean 55.72 61.88
+Adv 31.89 36.23
+WRN34-10 Clean 56.45 62.38
+Adv 28.56 32.05
+D.6 Performance on Other Datasets
+D.6.1 Imagenette
+Results in Table 15.
+Table 15: Comparative performance of CLAT across various networks on Imagenette. Robustness is evaluated with white-box PGD-10 and Auto Attack.
+Model Method Clean (%) PGD-10 (%) AA (%)
+DN121 PGD-AT 83.40 61.78 51.23
+PGD-AT + CLAT 86.91 65.45 54.82
+WRN70-16 PGD-AT 90.20 67.96 58.91
+PGD-AT + CLAT 93.52 72.39 61.43
+RN50 PGD-AT 84.02 62.10 50.45
+PGD-AT + CLAT 87.11 64.89 54.31
+WRN34-10 PGD-AT 90.45 65.45 56.31
+PGD-AT + CLAT 93.21 69.82 60.04
+VGG19 PGD-AT 85.45 56.71 46.53
+PGD-AT + CLAT 89.72 59.45 51.22
+RN18 PGD-AT 83.01 60.04 49.01
+PGD-AT + CLAT 86.42 62.91 51.23
+D.6.2 ImageNet
+Results in Table 16.
+Table 16: Clean and Adversarial Accuracies (PGD-10) performance comparison on ImageNet.
+Model Method Clean Adv
+DN121 PGD-AT 63.25 32.56
+PGD-AT + CLAT 66.10 35.48
+RN50 PGD-AT 65.88 33.18
+PGD-AT + CLAT 67.12 36.91
+WRN34-10 PGD-AT 64.31 31.08
+PGD-AT + CLAT 66.12 33.59
+D.7 Robustness against Other Attacks
+We report that performance on varying strengths of PGD-10 in Appendix B.
+Table 17 highlights the robustness against stronger white-box attacks, including those not limited to $\ell_{\infty}$-bounded constraints.
+Table 17: Adversarial accuracies across various attacks on CIFAR-10, highlighting the difference between models trained with the baseline and those utilizing CLAT.
+Positive blue values indicate the improvement in performance achieved with CLAT.
+Attack Method DN121 (%) WRN34-10 (%)
+FAB PGD-AT 44.80 40.12
+PGD-AT + CLAT +3.70 +5.03
+StAdv PGD-AT 48.50 45.15
+PGD-AT + CLAT +0.91 +1.89
+Pixle PGD-AT 10.40 9.50
+PGD-AT + CLAT +2.21 +1.90
+PGD-$\ell_{2}$ ($\epsilon$ = 0.03) PGD-AT 61.79 60.25
+PGD-AT + CLAT +1.92 +1.45
+$\ell_{\infty}$ ($\epsilon$ = 0.03, 50 steps) PGD-AT 57.01 54.01
+PGD-AT + CLAT +1.92 +2.03
+$\ell_{\infty}$ ($\epsilon$ = 0.03, 100 steps) PGD-AT 56.89 53.12
+PGD-AT + CLAT +1.63 +1.89
+D.8 Additional Performance Comparisons to Baselines
+D.8.1 Stochastic Weight Averaging and Adversarial Weight Perturbation
+Table 18 shows results when CLAT is augmented with SWA (Hwang et al., 2020) and AWP (Wu et al., 2020a) techniques. Omitted values are not reported in the original work.
+Table 18: PGD-10 Adversarial Accuracies on CIFAR-10 and CIFAR-100 for PreAct RN-18 and WRN34-10 compared to baselines.
+Network Method CIFAR-10 CIFAR-100
+PreAct RN-18 AWP (Wu et al., 2020a) 55.39 30.71
+AWP + CLAT 58.41 33.97
+SWAAT (Hwang et al., 2020) 58.32 28.43
+SWAAT + CLAT 60.76 30.74
+WRN34-10 AWP 58.10 -
+AWP + CLAT 60.89 -
+SWAAT 61.45 31.97
+SWAAT + CLAT 63.82 34.55
+D.8.2 Data Augmentation Techniques
+Please see Table 19.
+Table 19: Adversarial accuracies on CIFAR-10 for PreAct RN-18 (using PGD-10) and WRN70-16 (using Auto Attack).
+Network Method Adv. Accuracy (%)
+PreAct RN-18 Data Aug (weak) (Li & Spratling, 2023) 50.34
+Data Aug (weak) + CLAT 54.01
+Data Aug (strong) 49.99
+Data Aug (strong) + CLAT 52.37
+WRN70-16 Wang et al. (2023) 70.69
+Wang et al. (2023) + CLAT 72.34
+D.9 Timing Comparisons
+For a fair comparison, we use the same GPU configuration and number of GPUs across all methods, as described in the methods section.
+For DN121, one epoch of PGD-AT takes 67 seconds, RiFT takes 56 seconds per epoch, CLAT takes 69 seconds per epoch, and AutoLoRA also takes 69 seconds per epoch.
+As performed in the original paper for optimal performance, RiFT models were adversarially trained for 110 epochs, each taking 67 seconds, followed by fine-tuning for an additional 10 epochs at 56 seconds per epoch.
+Consequently, the total training time for RiFT is 132 minutes, compared to 112 minutes for CLAT (70 epochs of adversarial training and 30 epochs of fine-tuning).
+D.10 Ablation: Effect of choosing largest critical indices/ most critical layers
+Please see Table 20.
+Table 20: Ablation study of CLAT layer choices on CIFAR-10: The columns present Clean, PGD-10, and Auto Attack (AA) evaluation accuracies for models trained with CLAT, where layers with the largest and lowest values are selected.
+The optimal number of layers per network was chosen for both approaches.
+All settings are consistent with the results in Table 1
+.
+Network Largest cidx Smallest cidx PGD-AT Clean PGD-10 AA clean PGD-10 AA clean PGD-10 AA DN121 81.03 60.60 49.91 80.50 59.25 48.81 80.05 58.15 47.56 RN50 83.78 59.54 49.45 82.30 57.01 47.10 81.38 56.35 46.22 RN18 83.89 55.37 42.86 82.56 54.01 40.91 81.46 53.63 40.48
+Appendix E Curvature-based weakness measurement
+The main paper defines feature weakness based on the feature variation under worst-case perturbation.
+However, due to the non-linear and non-convex nature of the neural network model, the weakness measurement may not be precise in more complicated model architectures with a mixed layer type, such as the Vision Transformer model.
+To this end, this section provides a more accurate curvature-based formulation on feature weakness, and shows how the proposed weakness metric is an approximation.
+We leave the utilization of the curvature-based weakness measurement on more complicated models as future work.
+Let’s start by considering the feature perturbation function $G_{i}(\cdot)$, which is defined at the output of layer $i$ on inputs close to a clean data point $x$:
+$G_{i}(z)=||F_{i}(z)-F_{i}(x)||_{2}^{2}.$ (7)
+The worst-case curvature of the function $G_{i}$ at the neighborhood of $z=x$ can be estimated following the formulation by Moosavi-Dezfooli et al. (2019) as
+$\nu_{i}(x)=\frac{\nabla G_{i}(x^{\prime})-\nabla G_{i}(x)}{||x^{\prime}-x||_{2}}=\frac{\nabla G_{i}(x^{\prime})}{||x^{\prime}-x||_{2}},$ (8)
+where $x^{\prime}$ is a worst-case perturbation (adversarial attack) maximizing $G_{i}(z)$ in the vicinity of $x$, and $\nabla G_{i}(x)=0$ by definition given it is a minimum.
+Following the observation in Moosavi-Dezfooli et al.
+(2019), a higher curvature indicates the feature to be more non-robust to adversarial examples.
+We can therefore use the curvature formulation $\nu_{i}(x)$ under a fixed perturbation budget $||x^{\prime}-x||_{p}\leq\epsilon$ to estimate the layer non-robustness, or weakness.
+As we use the feature weakness to derive both the layer criticality metric and the finetuning objective, having a gradient term in the layer weakness leads to the costly computation of higher-order gradients in the optimization.
+To avoid the high cost of computing higher-order gradients when optimizing with the curvature, the numerator in the curvature formulation can be further derived as
+$\nu_{i}(x)=\frac{\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}^{T}(F_{i}(x^{\prime})-F_{i}(x))}{||x^{\prime}-x||_{2}}.$ (9)
+In practice, it is also difficult to explicitly instantiate $\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}$ for a neural network.
+To this end, we simplify the formulation in Eq.
+9 by assuming $\frac{\partial F_{i}(x^{\prime})}{\partial x^{\prime}}$ as a uniform vector.
+This leads to our definition of the $\epsilon$-weakness of layer $i$’s feature as:
+$\mathcal{W}_{\epsilon}(F_{i})=\frac{1}{N_{i}}\mathbb{E}_{x}\left[\sup_{||\delta||_{p}\leq\epsilon}||F_{i}(x+\delta)-F_{i}(x)||_{2}\right],$ (10)
+where $N_{i}$ denotes the dimensionality of the output features at layer $i$, therefore normalizing the weakness measurement of layers with different output sizes.
+The weakness measurement is proportional to the curvature estimation in Eq.
+9.
+A higher weakness value indicates that the feature vector is more vulnerable to input perturbations.
+The functionality of cascading layers from 1 to $i$ affects the vulnerability of the hidden features, as described by this formulation.